How much soil containing 45% sand do you need to add to 1 ft3 of soil containing 15% sand in order to make a soil containing 35% sand?
Answer by ptaylor(2198)

Let x=amount of soil containing 45% sand that needs to be added
Now we know that the amount of pure sand in the soil that's added (0.45x) plus the amount of pure sand in the 1 cu ft(0.15*1) has to equal the amount of pure sand in the final mixture(0.35(1+x)). So our equation to solve is:
0.45x+0.15*1=0.35(1+x) get rid of parens
0.45x+0.15=0.35+0.35x subtract 0.35x and also 0.15 from each side of the equation
0.45x-0.35x+0.15-0.15=0.35x-0.35x+0.35-0.15 collect like terms
0.10x=0.20
x=2 cu ft-------------amount of soil containing 45% sand that needs to be added
CK
0.45*2+0.15*1=0.35(2+1)
0.90+0.15=1.05
1.05=1.05
Hope this helps---ptaylor